# Don’t Let It Hallucinate: Premise Verification via Retrieval-Augmented Logical Reasoning

### v1

@@ 1 Introduction
Prior research distinguishes between factuality hallucinations—where the output conflicts with known facts—and faithfulness hallucinations—where the response diverges from the provided context or user instructions (Snyder et al., 2024a). We focus on factuality hallucinations, particularly those driven by incorrect assumptions (false premises). These errors are common among LLM outputs (Huang et al., 2025; Snyder et al., 2024b), where Figure 1 presents an example of a question with a false premise and the resulting hallucination.
Figure 1: LLM experiences factuality hallucination when faced with a false premise question, where both entities (The Lord of the Rings: The Return of the King, AACTA Award for Best Adapted Screenplay) exist but are not correctly aligned.
@@
To address this challenge, we focus on preventing hallucinations rather than mitigating them post hoc. In our framework, outlined in Section 3, we first transform the user’s query into a logical form that highlights key entities or relations (§3.2). We then employ retrieval-augmented generation (RAG) to check the accuracy of these statements against a knowledge graph (§3.3). If contradictions are found, the query is flagged as containing a false premise (§3.1), prompting the model to correct or reject the assumption before formulating a final answer. This process, shown in Figure 2, ensures that the LLM does not rely on erroneous details during response generation. By informing the LLM about any detected false premise in advance (§3.4), we reduce the likelihood of hallucinations without requiring access to model logits or large-scale fine-tuning.
@@ 2 Related Works
False Premise. A False Premise Question (FPQ) is a question that contains incorrect facts that are not necessarily explicitly stated but might be mistakenly believed by the questioner (Yu et al., 2022; Kim et al., 2021). Recent studies (Yuan et al., 2024) have demonstrated that FPQs can induce factuality hallucination in LLMs, as these models often respond directly to FPQs without verifying their validity. Notably, existing prompting techniques like few-shot prompting (Brown et al., 2020) and Chain-of-Thought (Wei et al., 2023), tend to increase hallucinations. Conversely, directly prompting LLMs to detect false premises negatively impacts their performance on questions containing valid premises (Vu et al., 2023).
Knowledge Graph Fact Checking and Question Answering. In the related field of fact checking, Retrieval-Augmented Generation (RAG) approaches are utilized for checking whether data are true and accurate. Among different RAG tasks, knowledge graph-driven RAG has drawn attention due to its ability to effectively leverage structured knowledge. Recent works include 1) prompt based: (Pan et al., 2023) asks LLMs if evidence is sufficient. If not, they further ask LLMs questions critical for verifying a claim. (Sun et al., 2024) uses the LLM to retrieve relevant facts hop-by-hop. 2) graph based: (He et al., 2024) extracts subgraph from knowledge graph by formulating RAG over graph as Prize-Collecting Steiner Tree optimization problem. (Mavromatis & Karypis, 2024) uses graph neural network to reason over a dense subgraph and retrieve answer candidates for the given question. 3) training based: (Zheng et al., 2024) trains two encoders for retrieval: one for embedding the queries, and one for embedding the subgraph evidences. (Liu et al., 2024a) trains encoder to conduct the following retrieval and ranker processes. However, it requires that entity exists in the knowledge graph and uses results generated from prompts for training.
@@
Hallucination Mitigation. Sources of LLM hallucinations originate from different stages in the LLM life cycle (Zhang et al., 2023a), leading existing mitigation methods to target specific stages: 1) Pre-training: Enhancing factual reliability by emphasizing credible texts, either by up-sampling trustworthy documents (Touvron et al., 2023) or prepending factual sentences with topic prefixes (Lee et al., 2023). 2) Supervised Fine-tuning: Curating high-quality, instruction-oriented datasets (Chen et al., 2024; Cao et al., 2024) improves factual accuracy more effectively than fine-tuning on unfiltered data, and remains more feasible compared to extensive pre-training. 3) Reinforcement Learning from Human Feedback: Aligning closely with human preferences may inadvertently encourage hallucinations or biased outputs, especially when instructions surpass the model’s existing knowledge (Radhakrishnan et al., 2023; Wei et al., 2024). 4) Inference: Known as hallucination snowballing (Zhang et al., 2023b), LLMs occasionally magnify initial generation mistakes. Proposed inference-time solutions include new decoding strategies (Shi et al., 2023; Chuang et al., 2024), uncertainty analysis of model outputs (Xu & Ma, 2025; Liu et al., 2024b; Dhuliawala et al., 2023). However, these approaches either act post-hallucination or require access to model logits, thus being inefficient due to repeated prompting or limited to white-box LLM scenarios.
@@ 3 Methodology
Hallucinations in LLMs often stem from false premises in user queries. Instead of addressing hallucinations after they occur, we aim to prevent them by detecting (§3.1) and informing the presence of false premises to LLMs before response generation. Our proposed method achieves this through three key steps:
Logical Form Conversion. (§3.2) By converting the user query into a structured logical representation, we extract its core meaning, making it easier to analyze its factual consistency. We demonstrate its effectiveness across various types of graph retrievers.
Structured Retrieval and Verification. (§3.3) Rather than relying solely on model-generated text, we retrieve external evidence to assess whether false premises exist in the query.
Factual Consistency Enforcement. (§3.4) The verified information is then incorporated into the LLM prompt, ensuring that the model generates responses aligned with factual data.
@@
3:Convert user query $q$ into logical representation $\mathcal{L}(q)$triangleright3.2)
@@
5:Initialize maximum similarity score $Sim_{max}\leftarrow-\infty$triangleright3.3)
@@
[…]
@@ 3.1 Problem Definition
$F(q)=\begin{cases}1,&\text{if }q\text{ conflicts with retrieved evidence }R(q,G),\\ 0,&\text{otherwise}.\end{cases}$ (1)
where $R$ represents the retrieval function that extracts relevant evidence from a knowledge graph $G$. The query $q$ is evaluated against $R(q,G)$, and if contradictions are found, $q$ is deemed to contain a false premise ($F(q)=1$); otherwise, it is considered valid ($F(q)=0$). In this study, the function $F$ is achieved by RAG using a retriever that leverages logical form and a knowledge graph.
@@ 3.2 Logical Form Extraction
Logical Form: A logical form is a structured representation of statements or queries expressed using symbolic logic. It provides a structured way to capture semantic relationships within sentences, enabling precise and systematic reasoning. Given a natural language sentence query $q$, its logical form $\mathcal{L}(q)$ can be represented as: $\mathcal{L}(q)=P(x_{1},x_{2},\dots,x_{n}),$ where $P$ denotes a predicate or relation, and $x_{1},x_{2},\dots,x_{n}$ are variables or constants representing entities or concepts extracted from $q$. For example, for the query: Was The Lord of the Rings: The Return of the King nominated for the Academy Award for Best Director? , its logical form is: Nominated("The Lord of the Rings: The Return of the King", "Academy Award for Best Director"). Here, "The Lord of the Rings: The Return of the King" and "Academy Award for Best Director" are entities, while Nominated represents the relation. GPT-4o-mini (OpenAI et al., 2024) is used for extracting logical forms from the queries. For an input query $q$, we first ask the LLM to generate the corresponding logical form $L(q)$. Then, we extract the source, relationship, and target from $L(q)$.
@@ 3.3 Retrieval
[…]
@@
After converting a user query $q$ into a logical form representation $L(q)$, the retriever encodes the logical form and the graph triples, searches through the knowledge graph $G$, and extracts the most relevant triple or subgraph, applying different selection criteria depending on the retriever used in our study. Therefore, formula 2 can be further formulated to:
$G^{*}=\text{Graph-Retriever}(\mathcal{L}(q),G)=\arg\max_{G\subseteq R(G)}p_{\theta}(G|\mathcal{L}(q),G)=\arg\max_{G\subseteq R(G)}\text{Sim}(\mathcal{L}(q),G).$
@@ 3.4 Hallucination Mitigation
$q=\begin{cases}q\text{ + `` Note: This question contains a false premise.''},&\text{if }F(q)=1,\\ q,&\text{otherwise}.\end{cases}$ (3)
where $q$ represents the modified query that explicitly flags the presence of a false premise when detected. Once the original query is updated, we evaluate LLM’s responses and measure the effectiveness of the ensuing hallucination mitigation.
@@ 4.1 Dataset
KG-FPQ (Zhu et al., 2024) is a dataset containing true and false premise questions that are constructed from the KoPL knowledge graph, a high-quality subset of Wikidata. In KG-FPQ, GPTs are used to generate TPQs based on true triplets from the knowledge graph, and FPQs are created by replacing the original object with the edited object from false triplets through string matching. We look into the discriminative task in the art domain in the dataset. In the discriminative tasks, LLMs are required to answer Yes-No questions with “Yes” or “No”. An example for FPQ in Yes-No format is that Is Hercules a cast member of ”The Lord of the Rings: the Return of the King?. More details of the dataset are in Appendix A.1.
@@ 4.2 Experiment Setting
Our approach mitigates hallucination through a two-step process: First, we detect false premises in the user query (§4.2.1, §4.2.2). Then, we use the result of false premise detection along with the original query when providing input to the LLM (§4.2.3).
@@ 4.2.1 False Premise Detection with Logical Form
In the false premise detection task, we look at different retrievers with and without the use of logical forms. Logical forms are used in 1) the retrieval stage, where the logical form $\mathcal{L}(q)$ is encoded to find the most relevant elements from knowledge $G$, and 2) the false premise detection stage, where the logical form is passed as input along with the retrieved evidence to LLM to determine whether the query contains false premise. The prompt detail is in Appendix A.3. We evaluate the use of logical forms in three configurations: 1) applying logical forms in both the retrieval stage and false premise detection stage, 2) using logical forms for retrieval and employing the original query for false premise detection, and 3) utilizing the original query for both stages.
4.2.2 False Premise Detection Baselines
@@
3) Non-parametric Retriever: G-retriever (He et al., 2024) uses Prize-Collecting Steiner Tree problem (PCST) algorithm for extracting relevant subgraph from the knowledge graph. It does not rely on a trained model with learnable parameters.
@@
We use GPT-3.5-Turbo as the LLM in the false premise detection task. These retrievers are included because they enable retrieval without task-specific fine-tuning, making them more adaptable across different domains. Unlike training-based retrievers, which require labeled data and extensive computation, non-parametric retriever uses structured knowledge, embedding-based retriever utilizes pre-trained encoders to transform queries and knowledge into a shared vector space for efficient retrieval, and LLM-based retrieval leverages pre-trained language models’ generalization abilities. This setup evaluates the impact of logical forms on retrieval efficiency without the overhead of model training.
@@
True Positives (TP%) 44.44 33.33 37.78 8.89
True Negatives (TN%) 73.33 80.00 100.00 93.33
False Positives (FP%) 26.67 20.00 0.00 6.67
False Negatives (FN%) 55.56 66.67 62.22 91.11
F1 Score (%) 59.70 48.78$\downarrow$ 54.84$\downarrow$ 16.16
Accuracy (%) 69.20 73.33 91.11 81.27
@@
True Positives (TP%) 44.44 37.78 37.78 8.89
True Negatives (TN%) 73.33 86.67 100.00 93.33
False Positives (FP%) 26.67 13.33 0.00 6.67
False Negatives (FN%) 55.56 62.22 62.22 91.11
F1 Score (%) 59.70 53.97$\downarrow$ 54.84$\downarrow$ 16.16
Accuracy (%) 69.20 79.69 91.11 81.27
@@
True Positives (TP%) 44.44 60.00 75.56 8.89
True Negatives (TN%) 73.33 86.67 80.00 93.33
False Positives (FP%) 26.67 13.33 20.00 6.67
False Negatives (FN%) 55.56 40.00 24.44 91.11
F1 Score (%) 59.70 73.97 84.47 16.16
Accuracy (%) 69.20 82.86 79.37 81.27
Table 1: Comparison of performance metrics across different retrieval methods using logical forms and/or original queries.
Figure 3: Comparison of performance metrics across different retrieval methods using logical forms and/or original queries.
4.2.3 Hallucination Mitigation Baselines

### v2

@@ 1 Introduction
Prior research distinguishes between factuality hallucinations, where the output conflicts with known facts, and faithfulness hallucinations, where the response diverges from the provided context or user instructions (Snyder et al., 2024a). We focus on factuality hallucinations, particularly those driven by incorrect assumptions (false premises). These errors are common among LLM outputs (Huang et al., 2025; Snyder et al., 2024b), where Fig. 1 presents an example of a question with a false premise and the resulting hallucination. These premise-driven errors are particularly insidious as they can appear factually sound while being fundamentally incorrect.
Figure 1: LLM experiences factuality hallucination when faced with a false premise question, where both entities shellfish allergy symptom and penicillin exist but are not correctly aligned. The LLM’s hallucinated response could delay life-saving treatment by incorrectly recommending antibiotics for allergic reactions.
@@
To address this challenge, we focus on preventing hallucinations rather than mitigating them post hoc. In our framework, we first transform the user’s query into a logical form that highlights key entities or relations. We then employ retrieval-augmented generation (RAG) to check the accuracy of these statements against a knowledge graph. If contradictions are found, the query is flagged as containing a false premise prompting the model to correct or reject the assumption before formulating a final answer. This process, shown in Fig. 2, ensures that the LLM does not rely on erroneous details during response generation. By informing the LLM about any detected false premise in advance, we reduce the likelihood of hallucinations without requiring access to model logits or large-scale fine-tuning.
@@ 2 Related Works
False Premise. A False Premise Question (FPQ) is a question containing incorrect facts that are not necessarily explicitly stated but might be mistakenly believed by the questioner (Yu et al., 2022; Kim et al., 2021). Recent studies (Yuan et al., 2024; Li et al., 2024; 2025c) have demonstrated that FPQs can induce factuality hallucination in LLMs, as they often respond directly to FPQs without verifying their validity. Notably, existing prompting techniques like few-shot prompting (Brown et al., 2020) and Chain-of-Thought (Wei et al., 2023), tend to increase hallucinations. Conversely, directly prompting LLMs to detect false premises degrades their performance on questions containing valid premises (Vu et al., 2023).
Logical Forms. Symbolic solvers and logical forms are applied to logical reasoning by grounding natural language in symbolic representations. The latest trend is integrating LLMs with symbolic solvers to enhance their performance Olausson et al. (2023); Pan et al. (2023a); Li et al. (2025b; 2023), where natural language is translated into symbolic logic forms and deterministic symbolic solvers are employed for inference, enabling more accurate logical problem-solving. Similarly, SymbCoT Xu et al. (2024) converts input text into symbolic formats such as first-order logic, generates reasoning plans through logical rule application, and verifies the reasoning process to ensure consistency. These methods demonstrate that incorporating symbolic improves the reliability and interpretability of LLM outputs, making them well-suited for tasks requiring logical consistency.
Knowledge Graph Fact Checking and Question Answering. In fact checking, RAG approaches verify data accuracy, with knowledge graph-driven RAG gaining attention for effectively leveraging structured knowledge. Recent works include: 1) prompt-based methods where (Pan et al., 2023b) evaluates evidence sufficiency and generates verification questions, and (Sun et al., 2024) performs hop-by-hop fact retrieval; 2) graph-based approaches where (He et al., 2024) formulates RAG as a Prize-Collecting Steiner Tree problem for subgraph extraction, and (Mavromatis and Karypis, 2024) uses graph neural networks for dense subgraph reasoning and answer retrieval; 3) training-based methods where (Zheng et al., 2024) develops dual encoders for query and subgraph evidence embedding, and (Liu et al., 2024a) trains encoders for retrieval and ranking processes, though requiring entity presence in the knowledge graph and relying on prompt-generated training data.
@@
Hallucination Mitigation. Sources of LLM hallucinations originate from different stages in the LLM life cycle (Zhang et al., 2023a; Li et al., 2026; Shawn et al., 2025), leading existing mitigation methods to target specific stages: 1) Pre-training: Enhancing factual reliability by emphasizing credible texts, either by up-sampling trustworthy documents (Touvron et al., 2023) or prepending factual sentences with topic prefixes (Lee et al., 2023). 2) Supervised Fine-tuning: Curating high-quality, instruction-oriented datasets (Chen et al., 2024; Cao et al., 2024) improves factual accuracy more effectively than fine-tuning on unfiltered data, and remains more feasible compared to extensive pre-training. 3) Reinforcement Learning from Human Feedback: Aligning closely with human preferences may inadvertently encourage hallucinations or biased outputs, especially when instructions surpass the model’s existing knowledge (Radhakrishnan et al., 2023; Wei et al., 2024). 4) Inference: Known as hallucination snowballing (Zhang et al., 2023b), LLMs occasionally magnify initial generation mistakes. Proposed inference-time solutions include new decoding strategies (Shi et al., 2023; Chuang et al., 2024), uncertainty analysis of model outputs (Xu and Ma, 2025; Liu et al., 2024b; Dhuliawala et al., 2023; Li et al., 2025a). However, these approaches either act post-hallucination or require access to model logits, thus being inefficient due to repeated prompting or limited to white-box LLM scenarios. We briefly discuss the comparison between our work and previous post-hoc hallucination mitigation method in Tab. 1 and detailed discussion can be found in Discussion and Appendix § A.4.
Method Training cost Number of tokens Training time Model agnostic Black-box Compatible
[…]
Ours Zero Original query + logical form Zero Yes Yes
Table 1: Comparison of training and compatibility between post-hoc method and our method.
@@ 3 Methodology
LLM hallucinations often stem from false premises in user queries. Instead of addressing hallucinations post-generation, we aim to prevent them by detecting and informing the presence of false premises to LLMs before response generation. Our proposed method achieves this through three key steps:
Logical Form Conversion. By converting the user query into a structured logical form representation, we extract its core meaning, making it easier to analyze its factual consistency. We demonstrate its effectiveness across various types of graph retrievers.
Structured Retrieval and Verification. Rather than relying solely on model-generated text, we retrieve external evidence to assess whether false premises exist in the query.
Factual Consistency Enforcement. The verified information is then incorporated into the LLM prompt, ensuring that the model generates responses aligned with factual data.
@@
3:Convert user query $q$ into logical representation $\mathcal{L}(q)$ $\triangleright$3.2)
@@
5:Initialize maximum similarity score $Sim_{max}\leftarrow$-$\infty$ $\triangleright$3.3)
@@
qq+" $\leftarrow q+\text{" Note: This question contains a false premise."}$
@@ 3.1 Problem Definition
$F(q)=\begin{cases}1,&\text{if }q\text{ conflicts with retrieved evidence }R(q,G),\\ 0,&\text{otherwise},\end{cases}$ (1)
where $R$ denotes the retrieval function that extracts relevant evidence from a knowledge graph $G$. The query $q$ is evaluated against $R(q,G)$, and if contradictions are found, $q$ is deemed to contain a false premise ($F(q)=1$); otherwise, it is considered valid ($F(q)=0$). In this study, the function $F$ is achieved by RAG using a retriever that leverages logical form and a knowledge graph.
@@ 3.2 Logical Form Extraction
Logical Form: A logical form is a structured representation of statements or queries expressed using symbolic logic. It provides a structured way to capture semantic relationships within sentences, enabling precise and systematic reasoning. Given a natural language sentence query $q$, its logical form $\mathcal{L}(q)$ can be represented as: $\mathcal{L}(q)=P(x_{1},x_{2},\dots,x_{n}),$ where $P$ denotes a predicate or relation, and $x_{1},x_{2},\dots,x_{n}$ are variables or constants representing entities or concepts extracted from $q$. For example, for the query:
Was The Lord of the Rings: The Return of the King nominated for the Academy Award for Best Director?
Its logical form is:
Nominated("The Lord of the Rings: The Return of the King", "Academy Award for Best Director")
Here, "The Lord of the Rings: The Return of the King" and "Academy Award for Best Director" are entities, while Nominated is the relation. GPT-4o-mini (OpenAI, 2024) is used for extracting logical forms from the queries. For an input query $q$, we first ask the LLM to generate the corresponding logical form $L(q)$. Then, we extract the source, relationship, and target from $L(q)$. The prompt for logical form conversion is included in Appendix § A.3.
To evaluate the quality of the logical form conversion, we ask two annotators to manually grade the generated logical forms using a three-point scale: 1 (do not match), 2 (partially match), and 3 (match). Each annotator grades 100 randomly sampled outputs. Across all 200 samples, the generated logical forms receive a score of 3 in all of the cases. More details are in Appendix §A.6.
@@ 3.3 Retrieval
$G^{*}$ $=\text{Graph-Retriever}(q,G)$
$=\arg\max_{G\subseteq R(G)}p_{\theta}(G\mid q,G)$
$=\arg\max_{G\subseteq R(G)}\text{Sim}(q,G),$ (2)
@@
After converting a user query $q$ into a logical form representation $L(q)$, the retriever encodes the logical form and the graph triples, searches through the knowledge graph $G$, and extracts the most relevant triple or subgraph, applying different selection criteria depending on the retriever used in our study. Therefore, formula 3.3 can be further formulated to:
$G^{*}$ $=\text{Graph-Retriever}(\mathcal{L}(q),G)$
$=\arg\max_{G\subseteq R(G)}p_{\theta}(G\mid\mathcal{L}(q),G)$
$=\arg\max_{G\subseteq R(G)}\text{Sim}(\mathcal{L}(q),G).$ (3)
@@ 3.4 Hallucination Mitigation
$q=\begin{cases}q+W,&\text{if }F(q)=1,\\ q,&\text{otherwise},\end{cases}$ (4)
where $W$ = "Note: This question contains a false premise.", and $q$ is the modified query that explicitly flags the presence of a false premise when detected. Once the original query is updated, we evaluate LLM’s responses and measure the effectiveness of the ensuing hallucination mitigation.
@@ 4.1 Dataset
KG-FPQ (Zhu et al., 2024) is a dataset containing true and false premise questions that are constructed from the KoPL knowledge graph, a high-quality subset of Wikidata. TPQs are generated from true triplets, while FPQs are created by replacing objects in false triplets via string matching. We evaluate the discriminative task in the art domain, where LLMs answer Yes-No questions (e.g., Is Hercules a cast member of ’The Lord of the Rings: the Return of the King’?). Dataset details are in Appendix §A.1.
The CREAK dataset (Onoe et al., 2021) is a benchmark for commonsense reasoning about entity knowledge. Unlike prior datasets focused on general physical or social scenarios, CREAK targets inferences that combine factual knowledge about specific entities with commonsense reasoning. It contains 13,000 human-authored English claims about entities labeled as true or false, along with a small contrast set. Each example requires understanding both factual attributes and commonsense implications. Dataset details are in Appendix §A.2.
The FEVER dataset Thorne et al. (2018) consists of natural-language claims labeled as Supported, Refuted, or Not Enough Information. Evidence is retrieved from Wikipedia, which functions as a natural-language knowledge base. Claims are created by modifying Wikipedia sentences and are subsequently verified independently, without access to their original sources. FEVER is widely adopted as a standard benchmark for fact verification and claim validation.
@@ 4.2 Experiment Setting
Our approach mitigates hallucination through a two-step process: First, we detect false premises in the user query. Then, we use the result of false premise detection along with the original query when providing input to the LLM. We use both the KG-FPQ dataset and the CREAK dataset for evaluating the premise detection task, and we use the KG-FPQ dataset for the hallucination mitigation task, since CREAK dataset contains statements, not questions.
@@ 4.2.1 False Premise Detection with Logical Form
In the false premise detection task, we look at different retrievers with and without the use of logical forms. Logical forms are used in 1) the retrieval stage, where the logical form $\mathcal{L}(q)$ is encoded to find the most relevant elements from knowledge graph $G$, and 2) the false premise detection stage, where the logical form is passed as input along with the retrieved evidence to LLM to determine whether the query contains false premise. The prompt detail is in Appendix A.3. We evaluate the use of logical forms in three configurations: 1) applying logical forms in both the retrieval stage and false premise detection stage, 2) using logical forms for retrieval and employing the original query for false premise detection, and 3) utilizing the original query for both stages. We further analyze the role of different components in the logical form through ablation studies in Appendix A.4.2.
4.2.2 False Premise Detection Methods
@@
3) Non-parametric Retriever: G-retriever (He et al., 2024) uses Prize-Collecting Steiner Tree algorithm for extracting relevant subgraph from the knowledge graph. It does not rely on a trained model with learnable parameters.
@@
5) SAC3 Zhang et al. (2024) is a hallucination detection approach that identifies hallucinations by assessing semantic-aware cross-check consistency, which involves generating semantically equivalent question perturbations and performing cross-model response consistency verification. We include SAC3 as baseline for the KG-FPQ dataset.
We use GPT-4o-mini as the LLM in the false premise detection task. Additional evaluations on other LLMs are provided in Appendix §A.4.4. These retrievers are included because they enable retrieval without task-specific fine-tuning, making them more adaptable across different domains. Unlike training-based retrievers, which require labeled data and extensive computation, non-parametric retriever uses structured knowledge, embedding-based retriever utilizes pre-trained encoders to transform queries and knowledge into a shared vector space for efficient retrieval, and LLM-based retrieval leverages pre-trained language models’ generalization abilities. This setup evaluates the impact of logical forms on retrieval efficiency without the overhead of model training.
Metrics. We evaluate the false premise detection task using TPR (true positive rate), TNR (true negative rate), FPR (false positive rate), FNR (false negative rate), F1 score, and accuracy of the model successfully identifying questions containing false premises or not. Here, a positive instance refers to a question that contains a false premise. Higher TPR indicates better detection of false premises.
@@
True Positives (TP%) 44.44 33.33 88.89 8.89
True Negatives (TN%) 73.33 80.00 86.67 93.33
False Positives (FP%) 26.67 20.00 13.33 6.67
False Negatives (FN%) 55.56 66.67 11.11 91.11
F1 Score (%) 59.70 48.78 87.89 16.16
Accuracy (%) 69.20 73.33 88.57 81.27
@@
True Positives (TP%) 44.44 37.78 82.22 8.89
True Negatives (TN%) 73.33 86.67 93.33 93.33
False Positives (FP%) 26.67 13.33 6.67 6.67
False Negatives (FN%) 55.56 62.22 17.78 91.11
F1 Score (%) 59.70 53.97 86.97 16.16
Accuracy (%) 69.20 79.69 83.81 81.27
@@
True Positives (TP%) 44.44 60.00 94.44 8.89
True Negatives (TN%) 73.33 86.67 99.05 93.33
False Positives (FP%) 26.67 13.33 0.95 6.67
False Negatives (FN%) 55.56 40.00 5.56 91.11
F1 Score (%) 59.70 73.97 97.12 16.16
Accuracy (%) 69.20 82.86 95.24 81.27
Table 2: KG-FPQ dataset: comparison of performance metrics across different retrieval methods using logical forms and/or original queries.
Direct Claim with RAG G-retriever GraphRAG/ToG
Original Query for Both Stages
True Positives (TP%) 72.5 62.3 24.6 89.9
True Negatives (TN%) 89.7 86.8 92.6 91.2
False Positives (FP%) 10.3 13.2 7.4 8.8
False Negatives (FN%) 27.5 37.7 75.4 10.1
F1 Score (%) 79.4 71.1 37.4 90.5
Accuracy (%) 81.0 74.5 58.4 90.5
True Positives (TP%) 72.5 76.8 36.2 89.9
True Negatives (TN%) 89.7 92.6 83.8 88.2
False Positives (FP%) 10.3 7.4 16.2 11.8
False Negatives (FN%) 27.5 23.2 63.8 10.1
F1 Score (%) 79.4 83.5 47.6 89.2
Accuracy (%) 81.0 84.7 59.9 89.1
Logical Form for Both Stages
True Positives (TP%) 72.5 88.4 92.8 92.8
True Negatives (TN%) 89.7 92.6 83.8 91.2
False Positives (FP%) 10.3 7.4 16.2 8.8
False Negatives (FN%) 27.5 11.6 7.2 7.2
F1 Score (%) 79.4 90.4 88.9 92.1
Accuracy (%) 81.0 90.5 88.3 92.0
Table 3: CREAK dataset: comparison of performance metrics across different retrieval methods using logical forms and/or original queries.
Direct WRAG (with RAG) GraphRAG/ToG
Orig LF-Retr LF-Both
True Positive (TP%) 18.8 96.1 94.1 94.1 88.2
True Negative (TN%) 98.5 61.2 83.7 83.7 83.7
False Positive (FP%) 1.5 38.8 16.3 16.3 16.3
False Negative (FN%) 81.2 3.9 5.9 5.9 11.8
F1 Score (%) 31.3 82.4 89.7 89.7 86.5
Accuracy (%) 58.4 79.0 89.0 89.0 86.0
Table 4: FEVER dataset: comparison of performance metrics across different retrieval methods using logical forms and/or original queries. Since g-retriever is a graph-based retriever and does not apply to non-graph data, we do not include it here.
Figure 3: KG-FPQ dataset: comparison of performance metrics across different retrieval methods using logical forms and/or original queries.
Method Accuracy Number of tokens Running time* Model agnostic Black-box Compatible
Contrastive 84.8 Original Query + Reasoning Step Context Retrieval Time Agnostic to No
Decoding (Length $\gg$ Logical Form) + 10.6s White Box Models
Our Method 89.5 Original Query + Logical Form Context Retrieval Time Yes Yes
+ 0.6s
Table 5: Comparison of performance and efficiency between contrastive decoding and our method on the KG-FPQ dataset. *Average running time of each query on NVIDIA RTX A6000 GPU using Llama-3.1-8B Instruct model. Both methods require context retrieval.
4.2.3 Hallucination Mitigation Methods